OpenStudy (anonymous):

Find the diagonal of a square whose sides are of the given measure. Given = 7√3

OpenStudy (anonymous):

@acxbox22 sorry I keep tagging you haha

OpenStudy (acxbox22):

okay so using the rule i gave you before the side is 7sqrt3 \[7\sqrt{3}\times \sqrt{2}=7\sqrt{6}\]

OpenStudy (anonymous):

oh ok haha thanks :)

OpenStudy (acxbox22):

np